# Sequence Compression Speeds Up Credit Assignment in Reinforcement Learning

### v1

@@ C.3 TDC
To see how these updates enforce Equation 7, assume that at time $t-1$ Equation 7 holds for $S_{t}$. Then at time $t$ we visit $S_{t}$ and wish to make the correction for the new visit. Because we maintain $\hat{V}(s^{\prime},s)=\hat{V}_{T(s^{\prime},s)}(s^{\prime})$, we can write
@@
In effect, the update subtracts the outdated estimate from the sum, replaces it with the newer estimate, updates the incremental average for the reward and corrects the denominator to account for the additional visit to $S_{t}$. Having defined $\tilde{V}_{t}$ we can now try to rewrite it in a form that looks more like a conventional TD update

### v2

@@ C.3 TDC
To see how these updates enforce Equation 7, assume that at time $t-1$ Equation 7 holds for $S_{t}$. Then, at time $t$, we visit $S_{t}$ and wish to make the correction for the new visit. Because we maintain $\hat{V}(s^{\prime},s)=\hat{V}_{T(s^{\prime},s)}(s^{\prime})$, we can write
@@
In effect, the update subtracts the outdated estimate from the sum, replaces it with the newer estimate, updates the incremental average for the reward, and corrects the denominator to account for the additional visit to $S_{t}$. Having defined $\tilde{V}_{t}$, we can now try to rewrite it in a form that looks more like a conventional TD update